# ICU-Sepsis: A Benchmark MDP Built from Real Medical Data

### v1

@@ 4.3 Constructing the ICU-Sepsis MDP
This effectively means that the MDP still only allows the admissible actions to be taken, since taking an inadmissible action is equivalent to choosing one of the admissible actions at random and transitioning accordingly. Therefore, all optimal policies for the restricted-action setting remain optimal, and all policies that take inadmissible actions in some states can be mapped to equivalent policies that only use admissible actions (by spreading the probability of inadmissible actions across the admissible actions). This design decision enables the use of RL algorithm implementations that are only compatible with MDPs that allow all actions in all states, without giving them access to inadmissible actions. An episode ends when the agent reaches the state corresponding to survival or death, after which it can be considered to always transition to $s_{\infty}$ with probability $1$ regardless of action taken. Therefore, the states corresponding to survival and death are called terminal states.
@@ 4.4 Computing the final parameters
The result is the transition function $T$ represented as a tensor of shape $|\mathcal{S}|\times|\mathcal{A}^{+}|\times|\mathcal{S}|$, and the reward and initial-state distribution functions vectors $R$ and $d_{0}$, respectively, both represented as vectors of length $|\mathcal{S}|$. While we have largely followed the work of Komorowski et al. (2018) in the formulation of the MDP, we have made two important changes. First, the discount factor $\gamma$ has been set to $1$ instead of $0.99$ to prioritize patient survival over treatment speed. Secondly, the transition threshold $\tau$ has been increased from $5$ to $20$ to enable more accurate estimation of transition probabilities. The effects of these changes are examined in Appendix B. The values for all the parameters are shown in Table 1.
@@ 5.1 Methodology
Hyperparameter tuning is performed using a random search, where each algorithm runs for $300,\!000$ episodes, averaged over eight random seeds for each hyperparameter setting, to maximize expected returns for the last 10% of the episodes. After approximating the best set of hyperparameters through the random search, each algorithm is run for $500,\!000$ episodes averaged over $1,\!000$ random seeds to ensure robustness in results. More details about the search and the final hyperparameter values are given in Appendix C. We say that an algorithm has converged if the average return over the last $1,\!000$ episodes are within $0.1\%$ of the average return over the last $10,\!000$ episodes. Since the goal is to find policies with a high expected return in the environment, the returns are not evaluated on a separate MDP built with held-out data, as ICU-Sepsis acts as the ground truth in this case, and generalization of policies to other environments or the real world is not being tested.
@@ Appendix A Reproducing the ICU-Sepsis Parameters
Appendix B Examining the Effect of the Transition Threshold
@@ B.1 Baseline results
B.2 Performance of various algorithms
@@
Therefore, with respect to the second and third research questions, we see that the expected returns and average episode lengths in the learned policies are unusually high, which do not reflect the numbers seen in the dataset. We posit that this might be happening because the agent has learned to exploit some rare actions in certain states which happened to result in good outcomes by chance. Since increasing the transition threshold removes such actions from the set of admissible actions, this behavior is not observed in the ICU-Sepsis MDP which has a higher transition threshold but is otherwise identical in the creation process to the Variant MDP. To further validate this theory, in Appendix B.3 we test the robustness of the three baseline policies: a random policy, the expert policy, and the optimal policy learned using value iteration for both ICU-Sepsis and the Variant.
@@
B.3 Effect of perturbations on the environments
@@
Appendix C Hyperparameters Search
@@
C.1 Random search setting
@@
C.2 Best set of approximated hyperparameters

### v2

@@ 4.3 Constructing the ICU-Sepsis MDP
This effectively means that the MDP still only allows the admissible actions to be taken, since taking an inadmissible action is equivalent to choosing one of the admissible actions at random and transitioning accordingly. Therefore, all optimal policies for the restricted-action setting remain optimal, and all policies that take inadmissible actions in some states can be mapped to equivalent policies that only use admissible actions (by spreading the probability of inadmissible actions across the admissible actions). This design decision enables the use of RL algorithm implementations that are only compatible with MDPs that allow all actions in all states, without giving them access to inadmissible actions. We discuss this decision of how inadmissible actions are handled in more detail in Appendix B.
An episode ends when the agent reaches the state corresponding to survival or death, after which it can be considered to always transition to $s_{\infty}$ with probability $1$ regardless of action taken. Therefore, the states corresponding to survival and death are called terminal states.
@@ 4.4 Computing the final parameters
The result is the transition function $T$ represented as a tensor of shape $|\mathcal{S}|\times|\mathcal{A}^{+}|\times|\mathcal{S}|$, and the reward and initial-state distribution functions vectors $R$ and $d_{0}$, respectively, both represented as vectors of length $|\mathcal{S}|$. While we have largely followed the work of Komorowski et al. (2018) in the formulation of the MDP, we have made two important changes. First, the discount factor $\gamma$ has been set to $1$ instead of $0.99$ to prioritize patient survival over treatment speed. Secondly, the transition threshold $\tau$ has been increased from $5$ to $20$ to enable more accurate estimation of transition probabilities. The effects of these changes are examined in Appendix C. The values for all the parameters are shown in Table 1.
@@ 5.1 Methodology
Hyperparameter tuning is performed using a random search, where each algorithm runs for $300,\!000$ episodes, averaged over eight random seeds for each hyperparameter setting, to maximize expected returns for the last 10% of the episodes. After approximating the best set of hyperparameters through the random search, each algorithm is run for $500,\!000$ episodes averaged over $1,\!000$ random seeds to ensure robustness in results. More details about the search and the final hyperparameter values are given in Appendix D. We say that an algorithm has converged if the average return over the last $1,\!000$ episodes are within $0.1\%$ of the average return over the last $10,\!000$ episodes. Since the goal is to find policies with a high expected return in the environment, the returns are not evaluated on a separate MDP built with held-out data, as ICU-Sepsis acts as the ground truth in this case, and generalization of policies to other environments or the real world is not being tested.
@@ Appendix A Reproducing the ICU-Sepsis Parameters
Appendix B Handling Inadmissible Actions
Recall from Section 4.3 that state-action pairs that occur $\tau$ or fewer times in the dataset (where $\tau$ is a hyperparameter) are inadmissible (that is, they cannot be taken) since the subsequent transition distribution is unknown. This presents a problem: many implementations of RL algorithms do not allow for different action sets in each state. This may present a challenge for researchers hoping to compare to baselines that lack this functionality. We therefore opted to design ICU-Sepsis to be compatible with two different perspectives.
In the first perspective, inadmissible actions cannot be taken in the states where they are inadmissible. A list of admissible actions for each state is provided in the extras/admissibleActions.txt file provided with the CSV files containing the dynamics, as well as under the admissible_actions key in the info dictionary provided by the Gym/Gymnasium API. Furthermore, the entries in the transition probability table and reward function that correspond to inadmissible state-action pairs can be ignored. This perspective is ideal, simulating a setting where inadmissible actions do not exist as options for the agent to consider.
In the second perspective, we ensure that ICU-Sepsis is compatible with software that requires all actions to be admissible in all states. A key goal under this perspective is to avoid the fabrication of artificial environment behavior if inadmissible actions are chosen by the agent (e.g., defining inadmissible actions to cause a transition to a state representing death to discourage the selection of inadmissible actions). Such artificial transitions are undesirable because they can alter various performance metrics (e.g., performance improvement and learning curve plots could be dominated by the speed with which agents learn not to take inadmissible actions, which is not the important part of the ICU-Sepsis simulation). Instead we view inadmissible actions as being truly inadmissible (they cannot be taken by the agent, and hypothetical transitions that result from these actions should not be considered). To achieve this, we consider how ICU-Sepsis could be designed so that when RL software selects inadmissible actions, these inadmissible actions are automatically modified to correspond to admissible actions, thereby ensuring that inadmissible actions are never chosen by the agent.
The key insight to enable this is the creation of a mapping from any policy that allows all actions to a corresponding policy that only selects admissible actions. Although the agent can learn and reason using a policy that can select all actions, the interactions with the environment (including evaluations of expected return) are equivalent to a corresponding policy that only selects admissible actions.
The most straightforward way to achieve these desired properties would be to define inadmissible actions to instead represent any one of the admissible actions. If there is only one inadmissible action, this essentially gives the agent two different ways to select one of the admissible actions. Critically, this does not mean that the inadmissible action is actually chosen and the simulated result is the outcome of the inadmissible action. Instead, this means that inadmissible actions can never be chosen and instead a redundant policy representation is used (a policy representation that allows for multiple ways of selecting one or more of the admissible actions).
However, this straightforward approach introduces a different issue: in standard RL implementations that require all actions to be allowed in all states, there may not be a mechanism to tell the agent that in some states two different actions actually correspond to a single action. When the agent selects one of two equivalent actions, it may not recognize that the outcome of the action provides information about both of the actions. That is, the agent will not necessarily generalize properly. This raises questions regarding the significance of the choice of which action inadmissible actions map to. To avoid these complexities, we opt to map inadmissible actions to a distribution over the admissible actions.
Specifically, we define inadmissible actions in a given state to be equivalent to a uniform random selection of the admissible actions in that state. This means that if an agent that requires all actions to be allowed in all states selects a inadmissible action, its policy is implicitly modified to uniformly randomly select an action from the admissible set of actions. This achieves the desired goals: it is realistic in that it completely disallows actions that it would be irresponsible to allow an RL agent to take (it does not provide hypothetical simulations of the outcomes of these uncertain and risky actions) and it avoids skewing performance metrics because the agent cannot achieve a significant initial increase in expected discounted return by simply learning to avoid inadmissible actions (a uniform random policy over all actions is now equivalent to a uniform random policy over the admissible actions). However, it is worth noting the limitation that agents selecting inadmissible actions may still fail to properly generalize, possibly resulting in slower learning than agents that properly handle admissible action sets.
Appendix C Examining the Effect of the Transition Threshold
@@ C.1 Baseline results
C.2 Performance of various algorithms
@@
Therefore, with respect to the second and third research questions, we see that the expected returns and average episode lengths in the learned policies are unusually high, which do not reflect the numbers seen in the dataset. We posit that this might be happening because the agent has learned to exploit some rare actions in certain states which happened to result in good outcomes by chance. Since increasing the transition threshold removes such actions from the set of admissible actions, this behavior is not observed in the ICU-Sepsis MDP which has a higher transition threshold but is otherwise identical in the creation process to the Variant MDP. To further validate this theory, in Appendix C.3 we test the robustness of the three baseline policies: a random policy, the expert policy, and the optimal policy learned using value iteration for both ICU-Sepsis and the Variant.
@@
C.3 Effect of perturbations on the environments
@@
Appendix D Hyperparameters Search
@@
D.1 Random search setting
@@
D.2 Best set of approximated hyperparameters